OpenStudy (anonymous):

Simplify (5m2n-3)2. Select one: 5m^4n^6 25m^4/n^6 25m^2n^-6 5m^2/n^-6

OpenStudy (solomonzelman):

\[(5m^2n^{-3})^2~~~~~~~~~~~~~~~~THIS~?\]

OpenStudy (anonymous):

Yeah

OpenStudy (solomonzelman):

\[\huge\color{blue}{ a^{-b}= \frac{1}{a^b} } \]\[\huge\color{blue}{ s^d \times s^x=s^{d+x} } \]

OpenStudy (isaiah.feynman):

One last rule you need for this problem \[(a^{m})^{n}= a ^{mn}\]

OpenStudy (anonymous):

pl

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

5m^2/n^-6?